• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Deur A.

    Log entry time 16:03:24 on October 5,1998

    Entry number 2447

    keyword=shift summary

    shift summary (day shift 10/05/98)

    9:05 Mott measurement done yesterday: 72.1 % +/-0.6 beam pol.
    10:15 coda crashed (end-of-run crash)
    10:58 NMR measurement 30% target pol.
    11:00 change configuration to perpendicular runs
    11:21 launch dataspy: correlation between the coda deatime peaks and the hits on the plane 3 of hadron arm VDC. correlation seen eagh time we get a peak
    12:50 beam problems (injector)
    12:30 beam back
    12:58 end of the perpendicular runs @ P=0.782 GeV/c tot # of events: 7.9 M
    NMR measurement: 28% target polarization
    13:12 Begin longitudinal run @ Ep=0.666 GeV, E=1,719 GeV, theta=15.5
    13:59 Beam down (Moller for Hall A and B)
    14:48 Beam back
    15:10, coda very slow, have to reboot the crates and adaqs2
    15:27 coda back.